OpenStudy (igster1112):

Shawna reduced the size of a rectangle to a height of 2 in. What is the new width if it was originally 24 in wide and 12 in tall?

OpenStudy (studygurl14):

\(\large\frac{12}{2}=\frac{24}{w}\rightarrow12w=24(2)\rightarrow 12w = 48\rightarrow width=4\)